5raised to 5 raised to5 .......(24 times ) divide by 24 will have ___________ fractional value left .

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question
Given5raised to 5 raised to5 .......(24 times ) divide by 24
This can be written as (55)24=(5120)/24
in 5120last digit is always 5
The 10’s digit is 2 as base 10’s digit is even and also power is even
So 5120/24=k+(25/24)
so it will leave a remainder of 1